# Bifurcations and canards in two coupled FitzHugh–Nagumo equations July 25, 2025

### v1

@@ 4. Synchrony
Figure 4. Several singular solutions of (2) starting in region $A_{2}$ shown in two different projections, parameters $b=0.5$, $c=0$ and $k=1$. Trajectories starting close to the fold line are attracted to the synchrony plane; other trajectories go to $\operatorname{Fix}(\delta)$. Conventions: trajectories with initial conditions at the brown and gray dots, slow part in brown/gray solid lines, fast part dotted lines, equilibria white dots, fold lines (red/black), double fold points red dots, intersection with synchrony/anti-synchrony planes (solid/dotted lines, respectively. (A) - Projection of $C_{0}$ into the subspace $\left(x_{1},x_{2},y_{1}\right)$ (purple) and trajectories. (B) - Same trajectories as in (A) projected into the plane $\left(x_{1},x_{2}\right)$.
Figure 5. Two singular solutions of (2) starting in $\operatorname{Fix}(\gamma)\cap A_{2}$ and in $\operatorname{Fix}(\delta)\cap A_{3}$ shown in two different projections, parameters $b=0.1$, $c=0$ and $k=0.1$. Two distinct stable solutions - bistability. Conventions: trajectories with initial conditions at the brown and gray dots, slow part in brown/gray solid lines, fast part dotted lines, equilibria g white dots, fold lines (red/black), double fold points red dots, intersection with synchrony/anti-synchrony planes (solid/dotted lines), respectively. (A) - Projection of $C_{0}$ into the subspace $\left(x_{1},x_{2},y_{1}\right)$ (purple) and trajectories. (B) - Same trajectories as in (A) projected into the plane $\left(x_{1},x_{2}\right)$.
It remains to see whether the intersection of the the synchrony plane with the attracting part, $A_{\varepsilon}$, of $C_{\varepsilon}$ is also attracting.
Proposition 4.2.
For small $\varepsilon>0$ the intersection $\operatorname{Fix}(\gamma)\cap A_{\varepsilon}$ is locally flow-invariant under (2). It is normally hyperbolic and locally attracting if $b+2k\geq 0$.
@@
In the proof of Proposition 4.1 it is shown that if $b+2k\geq 0$ then, at any point $Z\in\operatorname{Fix}(\gamma)\cap A$, the component $N_{\gamma}(Z)$ of $D\hat{V}(Z)$ perpendicular to $\operatorname{Fix}(\gamma)$ satisfies $\det\left(N_{\gamma}\right)(Z)>1$ and $\operatorname{Tr}\left(N_{\gamma}\right)(Z)<0$. From Corollary 3.4 we know that $A_{\varepsilon}$ is ${\mathcal{O}}(\varepsilon)$ close to $A$. Therefore, for small enough $\varepsilon$, if $\tilde{Z}\in\operatorname{Fix}(\gamma)\cap A_{\varepsilon}$, then $\det\left(N_{\gamma}\right)(\tilde{Z})>0$ and $\operatorname{Tr}\left(N_{\gamma}\right)(\tilde{Z})<0$ and the second statement follows. ∎
Proposition 4.2 is illustrated in Figures 23, 4 and 5. Figures 4 and 5 illustrate the meaning of locally attracting: the synchrony plane attracts trajectories in an open set around it, but not all trajectories. In the example of the figures $c=0$, hence (2) has additional symmetry and this is explored in the next result.
Proposition 4.3.
@@
The intersection $\operatorname{Fix}(\delta)\cap A$ is flow-invariant and it is normally hyperbolic if and only if $b\geq 0$ and in this case it is locally attracting. Moreover, for small $\varepsilon\geq 0$ the intersection $\operatorname{Fix}(\delta)\cap A_{\varepsilon}$ is locally flow-invariant, normally hyperbolic and locally attracting if $b\geq 0$.
Proof.
The plane $\{(x,-x,y,-y)\}=\operatorname{Fix}(\delta)$ is flow-invariant since it is fixed by the additional symmetry $\delta$. It remains to check that it is normally hyperbolic and to see where it is attracting. In the coordinates $z_{j}$ used in the proof of Proposition 4.1 the subspace $\operatorname{Fix}(\delta)$ is defined by the equalities $z_{1}=z_{2}=0$, $z_{3}=2x$ and $z_{4}=2y$. The expressions for the two matrices $\left.D{\widehat{V}}(z_{1},z_{2},z_{3},z_{4})\right|_{(2x,2y,0,0)}$ and $\left.D{\widehat{V}}(z_{1},z_{2},z_{3},z_{4})\right|_{(0,0,2x,2y)}$ are the same. Hence, the component $N_{\delta}$ of $D{\widehat{V}}$ in the directions perpendicular to $\operatorname{Fix}(\delta)$ is given by:
$N_{\delta}=\begin{pmatrix}\varphi^{\prime}(x)&-1\\ 1&-b\end{pmatrix}$
@@
$\det(N_{\delta})=-b\varphi^{\prime}(x)+1\quad\text{and}\quad\operatorname{Tr}(N_{\delta})=\varphi^{\prime}(x)-b.$
Using the arguments of the proof of Proposition 4.1 it follows that $\operatorname{Fix}(\delta)\cap A$ is normally hyperbolic if and only if $b\geq 0$ and in this case it is locally attracting. The last statement follows from the same arguments as Proposition 4.2. ∎
When $c=0$, solutions of (2) that lie in the invariant plane $\operatorname{Fix}(\delta)$ satisfy $(x_{2}(t),y_{2}(t))=-(x_{1}(t),y_{1}(t))$. We say they correspond to two cells with antisynchrony.
@@
If $b>0$ then for sufficiently small $c\neq 0$ and for every $r>0$ there is a $C^{r}$ locally flow-invariant locally attracting manifold for (2) close to $\operatorname{Fix}(\delta)\cap A_{\varepsilon}$. If moreover $b+2k>0$ there is bistability in the sense that the two locally invariant manifolds $\operatorname{Fix}(\delta)\cap A_{\varepsilon}$ and $\operatorname{Fix}(\gamma)\cap A_{\varepsilon}$ are locally attracting.
@@
The result follows from the normal hyperbolicity of Propositions 4.1 and 4.3 and Proposition 4.2 and from the persistence of normally hyperbolic locally invariant manifolds proved by Fenichel in [6] (see also [19, Section 2.2]). ∎
Similarly, from the persistence of normally hyperbolic invariant manifolds we obtain:
@@
Then for sufficiently small $|b-\tilde{b}|$, $|c-\tilde{c}|$ and $|k-\tilde{k}|$ if $b+2k>0$ and $\tilde{b}+2\tilde{k}>0$, there is an attracting normally hyperbolic locally invariant manifold for the perturbed system close to $\operatorname{Fix}(\gamma)\cap A_{\varepsilon}$. In other words, the perturbed system has stable solutions with the two cells approximately synchronised. Moreover if $b>0$ and $\tilde{b}>0$ then, for sufficiently small $c$ and $\tilde{c}$, the perturbed system has an attracting normally hyperbolic locally invariant manifold close to $\operatorname{Fix}(\delta)\cap A_{\varepsilon}$, and hence it has bistability.
@@ 5. Canards and mixed-mode oscillations
As remarked in Section 3 trajectories called canards may follow the unstable part of the slow manifold for a considerable amount of time. In geometric terms a canard solution corresponds to the intersection of an attracting and a repelling slow manifold near a non-hyperbolic point of $\Sigma$, the set of fold lines defined in (3). In this section, we establish some hypotheses ensuring the existence of canards for (2).
@@
Szmolyan & Wechselberger [25] and Krupa et al. [17] have established that canards appear around some points where, after a rescaling that we will describe below, the slow equations have an equilibrium at a fold point.
@@
A point $X^{*}=(x_{1}^{*},x_{2}^{*},y_{1}^{*},y_{2}^{*})$ lies on the fold line $\Sigma$ if and only if either $\varphi^{\prime}(x_{1}^{*})=0$ or $\varphi^{\prime}(x_{2}^{*})=0$. Without loss of generality we take $\varphi^{\prime}(x_{1}^{*})\neq 0$ and $\varphi^{\prime}(x_{2}^{*})=0$, hence $x_{2}^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$, the case $\varphi^{\prime}(x_{1}^{*})=0$ and $\varphi^{\prime}(x_{2}^{*})\neq 0$ being identical, due to the symmetry. The case $\varphi^{\prime}(x_{1}^{*})=0=\varphi^{\prime}(x_{2}^{*})$ is treated separately at 5.2 below.
Since $\varphi^{\prime}(x_{2}^{*})=0$, the equation obtained from $\dot{y}_{2}$ in (6) yields no dynamical information at $x_{2}=x_{2}^{*}$. We overcome this by a time rescaling of $\tau=t/\varphi^{\prime}(x_{2})$, that is singular at $x_{2}^{*}$, as mentioned above. Writing $dx_{i}/d\tau=x_{i}^{\prime}$ the equations (6) transform into:
@@
The choice $x_{2}^{*}=2\sigma/\sqrt{3}$ imposes conditions on $x_{1}^{*}$ as shown in the next result.
@@
For every $k\neq 0$ and for any choice of the parameters $b$ and $c$ there is an equilibrium $(x_{1}^{*},x_{2}^{*})$ of (7) with $x_{2}^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$ and $\left|x_{1}^{*}\right|>\dfrac{2}{\sqrt{3}}$. Therefore the point $X^{*}=(x_{1}^{*},x_{2}^{*},\varphi(x_{1}^{*}),\varphi(x_{2}^{*}))$ lies at the boundary of the attracting part $A$ of $C_{0}$.
@@
At $x_{2}^{*}=2\sigma/\sqrt{3}$ we have $\varphi^{\prime}(x_{2}^{*})=0$ so $x_{1}^{\prime}=0$. Therefore $\left(x_{1}^{*},2\sigma/\sqrt{3}\right)$, $\sigma=\pm 1$, is an equilibrium of (7) if and only if $x_{2}^{\prime}=0$ i.e. when
@@
and the result follows immediately since for $|x|>\dfrac{2}{\sqrt{3}}$ the function $\varphi(x)$ takes all the values in ${\mathbf{R}}$. ∎
@@
The time rescaling that we have used allows us to gain enough hyperbolicity to obtain a complete analysis by standard methods from dynamical systems theory. We call the point $(x_{1}^{*},x_{2}^{*})$ obtained in Lemma 5.2 a folded equilibrium of (7).
@@
For $k\neq 0$, a folded equilibrium $(x_{1}^{*},2\sigma/\sqrt{3})$, with $\sigma=\pm 1$, of the equation (7) is either a saddle or an unstable node or an unstable focus or a saddle-node. For an open set of the parameters $(k,b,c)$ it is either a saddle or an unstable node.
@@
For $H(x_{1},x_{2})=\left(H_{1}(x_{1},x_{2}),H_{2}(x_{1},x_{2})\right)$ the derivative $DH(x_{1},x_{2})$ at a folded equilibrium $(x_{1}^{*},x_{2}^{*})$, $x_{2}^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$, of (7) is given by

### v3

@@ 4. Synchrony and bistability
Figure 5. Illustration of Theorems 4.2 and 4.6. Several singular solutions of (2) starting in region $A_{2}$ shown in two different projections, parameters $b=0.5$, $c=0$ and $k=1$. Trajectories starting close to the fold line are attracted to the synchrony plane; other trajectories go to $\operatorname{Fix}(\delta)$. Conventions: trajectories with initial conditions at the brown and gray dots, slow part in brown/gray solid lines, fast part dotted lines, equilibria white dots, fold lines (red/black), double fold points red dots, intersection with synchrony/anti-synchrony planes (solid/dotted lines), respectively. (A) - Projection of $C_{0}$ into the subspace $\left(x_{1},x_{2},y_{1}\right)$ (purple) and trajectories. (B) - Same trajectories as in (A) projected into the plane $\left(x_{1},x_{2}\right)$.
Figure 6. Illustration of Theorems 4.2 and 4.4. Two singular solutions of (2) starting in $\operatorname{Fix}(\gamma)\cap A_{2}$ and in $\operatorname{Fix}(\delta)\cap A_{3}$ shown in two different projections, parameters $b=0.1$, $c=0$ and $k=0.1$. Two distinct stable solutions - bistability. Conventions: trajectories with initial conditions at the brown and gray dots, slow part in brown/gray solid lines, fast part dotted lines, equilibria white dots, fold lines (red/black), double fold points red dots, intersection with synchrony/anti-synchrony planes (solid/dotted lines), respectively. (A) - Projection of $C_{0}$ into the subspace $\left(x_{1},x_{2},y_{1}\right)$ (purple) and trajectories. (B) - Same trajectories as in (A) projected into the plane $\left(x_{1},x_{2}\right)$.
It remains to see whether the intersection of the synchrony plane with the attracting part, $A_{\varepsilon}$, of $C_{\varepsilon}$ is also attracting.
Lemma 4.4.
For small $\varepsilon>0$ the intersection $\operatorname{Fix}(\gamma)\cap A_{\varepsilon}$ is locally flow-invariant under (2). It is normally hyperbolic and locally attracting if $k>-b/2$.
@@
In the proof of Lemma 4.3 it is shown that if $k\geq-b/2$ then, at any point $Z\in\operatorname{Fix}(\gamma)\cap A$, the component $N_{\gamma}(x)$ of $D\hat{V}(Z)$ perpendicular to $\operatorname{Fix}(\gamma)$ satisfies $\det\left(N_{\gamma}(x)\right)>0$ and $\operatorname{tr}\left(N_{\gamma}(x)\right)<0$. From Corollary 3.3 we know that $A_{\varepsilon}$ is ${\mathcal{O}}(\varepsilon)$ close to the synchrony plane $A$, where $\varphi^{\prime}(x)<0$. Therefore, since $k\geq-b/2$, for small enough $\varepsilon$ and if $\tilde{Z}=(\tilde{x},\tilde{x},\tilde{y},\tilde{y})\in\operatorname{Fix}(\gamma)\cap A_{\varepsilon}$, then for sufficiently small $\varepsilon>0$ we have $\varphi(\tilde{x})<1/(b+2k)$. Therefore $\det\left(N_{\gamma}(\tilde{x})\right)>0$ and $\operatorname{tr}\left(N_{\gamma}(\tilde{x})\right)<0$ and the second statement follows. ∎
Equilibria in the synchrony plane satisfy both $y=\varphi(x)=4x-x^{3}$ and $x-b\varphi(x)-c=0$. Therefore in $\operatorname{Fix}(\gamma)$ there is always at least one equilibrium and there are at most three equilibria — details in Section 3 of [16]. By the Bézout Theorem there are at most 9 equilibria of (2) so there may be up to 3 symmetry related pairs of equilibria outside the synchrony plane. The next result is an example of the dynamics in the synchrony plane.
Corollary 4.5.
For small $\varepsilon>0$ if $b>1/4$ and $k>-1/8$ then for small values of $c$ the equations (2) have a synchronous asymptotically stable relaxation oscillation.
Proof.
The result follows from the fact that under these conditions on the parameters $b$ and $c$ the single FHN (1) has an asymptotically stable relaxation oscillation as shown in [16]. This follows by the Poincaré-Bendixson Theorem, because for small values of the parameter $c$ and $b>1/4$ there is only one equilibrium of (1) and it lies on the repelling part of the critical manifold.
The dynamics in the synchrony plane is that of a single FHN, hence (2) has a synchronous relaxation oscillation that, in the restriction to the synchrony plane, is asymptotically stable. If $k>-1/8$ then Theorem 4.2 holds, so the intersection of the synchrony plane with $A_{\varepsilon}$ is attracting, therefore the relaxation oscillation is asymptotically stable. ∎
Theorem 4.2 is illustrated by singular solutions shown in Figures 34, 5 and 6. In the example of Figure 4 we have $b=c=0$, so this is the interaction of two van der Pol equations (see, for instance [25, Section 1.3]). Figures 5 and 6 illustrate the meaning of locally attracting: the synchrony plane attracts trajectories in an open set around it, but not all trajectories. In the examples of the figures we have $c=0$, hence (2) has additional symmetry and this is explored in the next result.
Theorem 4.6.
If $c=0$ then the antisynchrony plane ${\mathcal{A}S}$ is flow-invariant under (2). For small $\varepsilon>0$ if $c=0$ then the intersection ${\mathcal{A}S}\cap A_{\varepsilon}$ of the antisynchrony plane with the attracting part of the slow manifold $C_{\varepsilon}$, $A_{\varepsilon}$, is locally flow-invariant under (2). It is normally hyperbolic and locally attracting if $b>0$.
Proof.
@@
As in the proof of Theorem 4.2 the plane $\{(x,-x,y,-y)\quad x,y\in{\mathbf{R}}\}=\operatorname{Fix}(\delta)={\mathcal{A}S}$ of points fixed by $\delta$ is flow-invariant. It remains to check that it is normally hyperbolic and to see where it is attracting. In the coordinates $z_{j}$ used in the proof of Lemma 4.3 the subspace $\operatorname{Fix}(\delta)$ is defined by the equalities $z_{1}=z_{2}=0$, $z_{3}=2x$ and $z_{4}=2y$. The expressions for the two matrices $\left.D{\widehat{V}}(z_{1},z_{2},z_{3},z_{4})\right|_{(2x,2y,0,0)}$ and $\left.D{\widehat{V}}(z_{1},z_{2},z_{3},z_{4})\right|_{(0,0,2x,2y)}$ are the same. Hence, the component $N_{\delta}(x)$ of $D{\widehat{V}}$ in the directions perpendicular to $\operatorname{Fix}(\delta)$ is given by:
$N_{\delta}(x)=\begin{pmatrix}\varphi^{\prime}(x)/\varepsilon&-1/\varepsilon\\ 1&-b\end{pmatrix}$
@@
$\det(N_{\delta}(x))=-\frac{1}{\varepsilon}(b\varphi^{\prime}(x)-1)\quad\text{and}\quad\operatorname{tr}(N_{\delta}(x))=\frac{\varphi^{\prime}(x)}{\varepsilon}-b.$
Using the arguments of the proof of Lemma 4.3 it follows that $\operatorname{Fix}(\delta)\cap A$ is normally hyperbolic if and only if $b\geq 0$ and in this case it is locally attracting. The last statement follows from the same arguments as Lemma 4.4. ∎
@@
If $b>0$ then for sufficiently small $c\neq 0$ and for every $r>0$ there is a $C^{r}$ locally flow-invariant locally attracting manifold for (2) close to $\operatorname{Fix}(\delta)\cap A_{\varepsilon}$. If moreover $k>-b/2$, then there is bistability in the sense that this manifold and the locally invariant manifold $\operatorname{Fix}(\gamma)\cap A_{\varepsilon}$ are simultaneously locally attracting.
@@
The result follows from the normal hyperbolicity established in Theorems 4.2 and 4.6 and from the persistence of normally hyperbolic locally invariant manifolds proved by Fenichel in [12] (see also [25, Section 2.2]). ∎
Similarly, from the persistence of normally hyperbolic invariant manifolds we obtain persistence of the dynamics with respect to the parameters. For this we need a definition.
Definition 4.8.
Let $\xi(t)=\left(x_{1}(t),x_{2}(t),y_{1}(t),y_{2}(t)\right)$ be a solution of a system of two coupled equations
$\left\{\begin{array}[]{rcl}\varepsilon\dot{x}_{i}&=&f_{i}(x_{1},x_{2},y_{1},y_{2})\\ \dot{y}_{i}&=&g_{i}(x_{1},x_{2},y_{1},y_{2})\end{array}\right.\qquad i=1,2.$
Then we define:
  • $\xi(t)$ is in approximate synchrony with precision $\Delta>0$ if there is a $t_{0}\in{\mathbf{R}}$ such that for all $t>t_{0}$ we have $\left|x_{1}(t)-x_{2}(t)\right|<\Delta$ and $\left|y_{1}(t)-y_{2}(t)\right|<\Delta$;
  • $\xi(t)$ is in approximate antisynchrony with precision $\Delta>0$ if there is a $t_{0}\in{\mathbf{R}}$ such that for all $t>t_{0}$ we have $\left|x_{1}(t)+x_{2}(t)\right|<\Delta$ and $\left|y_{1}(t)+y_{2}(t)\right|<\Delta$.
For the original system (2) we obtain from Corollary 4.7 a result analogous to Corollary 4.5.
Corollary 4.9.
For small $\varepsilon>0$ if $b>1/4$ then for small values of $c$ the equations (2) have an asymptotically stable relaxation oscillation in approximate antisynchrony with precision $\Delta={\mathcal{O}}(c)$. If moreover $k>-1/8$ then the attracting relaxation oscillations in synchrony and antisynchrony (in ${\mathcal{S}}$ and ${\mathcal{A}S}$) coexist.
If the two coupled FHN have slightly different parameter values, then synchrony and antisynchrony still hod in the approximate sense above, as in the next result.
@@
Then if $k>-b/2$ and $\tilde{k}>-\tilde{b}/2$, for small $\varepsilon>0$ and for sufficiently small $|b-\tilde{b}|$, $|c-\tilde{c}|$, $|k-\tilde{k}|$ the perturbed system has stable solutions with approximate synchrony with precision $\Delta={\mathcal{O}}\left(|b-\tilde{b}|,|c-\tilde{c}|,|k-\tilde{k}|\right)$.
If moreover $b>0$ and $\tilde{b}>0$ then, for sufficiently small $c$ and $\tilde{c}$, the perturbed system has stable solutions with approximate antisynchrony with precision $\Delta={\mathcal{O}}\left(|b-\tilde{b}|,|c-\tilde{c}|,|k-\tilde{k}|\right)$, and hence it has bistability.
Proof.
From the normal hyperbolicity established in Theorems 4.2 and 4.6 and from Fenichel’s theorem [12] it follows that for sufficiently small $|b-\tilde{b}|$, $|c-\tilde{c}|$ and $|k-\tilde{k}|$ if $k>-b/2$ and $\tilde{k}>-\tilde{b}/2$ there is an attracting normally hyperbolic locally invariant manifold for the perturbed system close to $\operatorname{Fix}(\gamma)\cap A_{\varepsilon}$. Also if $b>0$ and $\tilde{b}>0$ then, for sufficiently small $c$ and $\tilde{c}$, the perturbed system has an attracting normally hyperbolic locally invariant manifold close to $\operatorname{Fix}(\delta)\cap A_{\varepsilon}$. ∎
@@ 5. Canards and mixed-mode oscillations
In Section 2 we stated that, in general, solutions of a fast-slow system evolve close to the slow manifold under the slow equations and jump out of it when they reach the fold points, where the slow manifold becomes unstable. However, as remarked in Section 3 there are special situations when trajectories called canards may follow the unstable part of the slow manifold for a considerable amount of time, as drawn in Figure 7. In geometric terms a canard solution corresponds to the intersection of an attracting and a repelling slow manifold near a non-hyperbolic point of $\Sigma$, the set of fold lines defined in (3). In this section, we establish some hypotheses ensuring the existence of canards for (2) and explore some of the consequences for the dynamics. A description of the dynamics of mixed-mode oscillations associated to canards in general can be found in Desroches et al. [7] and for the forced van der Pol equations in Guckenheimer [17].
Figure 7. Schematic representation of a canard: a trajectory that follows the unstable part of the slow manifold for some time, before jumping out of it. Here the critical manifold is the thick red curve, solid in the attracting part and dashed in the repelling part.
@@
Intuitively, we are looking at a trajectory that evolves near the attracting part of the critical manifold. Then it reaches a fold point, where the critical manifold changes from attracting to repelling with respect to the fast equations. A canard happens if around this point the fast equations not only have an equilibrium but are specially slow. In this way the slow equation dominates the dynamics and the trajectory continues to follow close to the critical manifold for some time, near its unstable part, like in Figure 7. In a more rigorous setting, Szmolyan and Wechselberger [34], and Krupa et al. [23] have established that, for one-dimensional and two-dimensional slow equations, canards appear around some points where, after a rescaling that we will describe below, the slow equations have an equilibrium at a fold point.
5.1. An open set of parameters where canards occur
@@
A point $X^{*}=(x_{1}^{*},x_{2}^{*},y_{1}^{*},y_{2}^{*})$ lies on the fold lines $\Sigma$ if and only if either $\varphi^{\prime}(x_{1}^{*})=0$ or $\varphi^{\prime}(x_{2}^{*})=0$. Without loss of generality we take $\varphi^{\prime}(x_{1}^{*})\neq 0$ and $\varphi^{\prime}(x_{2}^{*})=0$, hence $x_{2}^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$, the case $\varphi^{\prime}(x_{1}^{*})=0$ and $\varphi^{\prime}(x_{2}^{*})\neq 0$ being identical, due to the symmetry. The case $\varphi^{\prime}(x_{1}^{*})=0=\varphi^{\prime}(x_{2}^{*})$ is treated separately in Subsection 5.3 below.
Since $\varphi^{\prime}(x_{2}^{*})=0$, the equation obtained from $\dot{y}_{2}$ in (6) yields no dynamical information at $x_{2}=x_{2}^{*}$. We overcome this by a time rescaling of $\tau=t/\varphi^{\prime}(x_{2})$, that is singular at $x_{2}^{*}$ since $\varphi^{\prime}(x_{2}^{*})=0$, as mentioned above. Writing $dx_{i}/d\tau=x_{i}^{\prime}$ the equations (6) transform into:
@@
Let $H(x_{1},x_{2})=\left(H_{1}(x_{1},x_{2}),H_{2}(x_{1},x_{2})\right)$ for $(x_{1},x_{2})\in{\mathbf{R}}^{2}\backslash\{(x_{1},x_{2}):\varphi^{\prime}(x_{1})=0\}$ be the vector field corresponding to (7).
If the fold point $X^{*}$ is also an equilibrium of (7) then the fast equation is slow enough around it to yield a canard. However, we need to make sure that $X^{*}$ lies at the boundary of the of the attracting part $A$ of the critical manifold $C_{0}$. Given the choice $x_{2}^{*}=2\sigma/\sqrt{3}$ this imposes conditions on $x_{1}^{*}$, that we deal with in the next result.
@@
For every $k\neq 0$ and for any choice of the parameters $b$ and $c$ there is an equilibrium $(x_{1}^{*},x_{2}^{*})$ of (7) with $x_{2}^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$ and $\left|x_{1}^{*}\right|>\dfrac{2}{\sqrt{3}}$. Therefore the point $X^{*}=(x_{1}^{*},x_{2}^{*},\varphi(x_{1}^{*}),\varphi(x_{2}^{*}))$ lies at the boundary of the attracting part $A$ of the critical manifold $C_{0}$.
@@
Taking into account (6), at $x_{2}^{*}=2\sigma/\sqrt{3}$ we have $\varphi^{\prime}(x_{2}^{*})=0$ so $x_{1}^{\prime}=0$. Therefore $\left(x_{1}^{*},2\sigma/\sqrt{3}\right)$, $\sigma=\pm 1$, is an equilibrium of (7) if and only if $x_{2}^{\prime}=0$ i.e., when
@@
and the first part of the result follows immediately since for $|x|>2/\sqrt{3}$ the function $\varphi(x)$ takes all the values in ${\mathbf{R}}$. The second part is immediate taking into account that $\left|x_{1}^{*}\right|>2/\sqrt{3}$. ∎
@@
The time rescaling that we have used allows us to gain enough hyperbolicity to obtain a complete analysis by standard methods from dynamical systems theory. We call the point $(x_{1}^{*},x_{2}^{*})$ obtained in Lemma 5.2 a folded equilibrium of (7). For the sake of completeness, we are going to review a way to detect canards of (2) from folded equilibria, critical points where the slow manifold generically changes the stability. See also Guckenheimer & Haiduc [18] and Guckenheimer [19].
Definition 5.3 (Definition 2.1 of [34], adapted).
Solutions of the reduced problem (7) passing through a fold point from an attracting region to a repelling one are called singular canards.
In order to have a canard it is not enough to have a folded equilibrium—it is also necessary to have a trajectory that approaches it in positive time. This is the idea behind the next result.
Lemma 5.4 (Lemma 2.3 of [34], adapted).
A canard of (2) corresponds to, at least, one real negative eigenvalue of the matrix $DH$ at a folded equilibrium.
The proof of Lemma 5.4 uses blow-up of singularities and foliation by center manifolds as main techniques. The blowing up makes the fine structure of the singular perturbation problem visible and then the full problem can be analysed by standard methods from dynamical systems theory. We return now to our original problem.
@@
For $k\neq 0$, a folded equilibrium $(x_{1}^{*},2\sigma/\sqrt{3})$, with $\sigma=\pm 1$, of the equation (7) is either a saddle or an unstable node or an unstable focus or a saddle-node. There exists a non-empty open set of the parameters $(k,b,c)\in{\mathbf{R}}^{3}$ for which it is either a saddle or an unstable node.
@@
The derivative $DH(x_{1},x_{2})$ at a folded equilibrium $(x_{1}^{*},x_{2}^{*})$, $x_{2}^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$ is given by